# Thompson sampling MAB multi-armed bandit

### v2

@@ II-C Lower bound on regret
Consider the MAB problem where the objective is to optimize the variance of the observed rewards. Let $\pi$ be an $\alpha$-consistent policy and $\Theta=\left\{\left\{\theta_{i}\right\}_{i=1}^{K}\colon\theta_{i}\in\mathcal{U}\right\}$ be the distribution model. Under Assumption 1, for any constant $C_{1}<1-\alpha$, $10<C_{2}<$ 1, with $f_{i}$ and $f_{*}$ as the reward distribution of the i-th arm and the optimal arm respectively, the model-specific regret satisfies,
$\liminf_{n\rightarrow\infty}\frac{\text{R}_{\texttt{VV}}(n)}{\log n}\geq\sum_{i=1,i\neq*}^{K}\frac{C_{1}C_{2}\delta_{i}}{I(f_{i},f_{*})}.$
@@
See Appendix VII-C. ∎
From Theorem 1 and Theorem 2, we see UCB-VV achieves an upper bound on the regret of $\mathcal{O}\log(n)$ and matches a lower bound of the same order up to constant factors, demonstrating that our proposed algorithm is asymptotically optimal.
@@ III-A Upper bound on error for BAI
Parameter: $\bar{V}_{i}=0$ Initialize: $A_{1}\leftarrow K$, $r\leftarrow 1$
@@ III-B Lower bound on error for BAI
$G^{k}$ =$=\begin{cases}\nu_{k}=\nu_{i}&i\neq k\\ \nu_{k}\sim U\left[\tilde{l}_{k},\tilde{u}_{k}\right]&i=k\end{cases}$
@@
$\mathbb{P}\left(\left|\frac{1}{n}\sum_{s=1}^{n}Z_{s}-\mathbb{E}[Z]\geq\epsilon\right|\right)\leq 2\exp\left(\frac{-2{n}\epsilon^{2}}{M^{2}}\right).$
@@ IV Sub-Gaussian Distributions and Sharpe Ratio
In this section, we extend our framework from bounded distributions to sub-Gaussian distributions. Specifically, we first derive a concentration inequality for the sample variance of sub-Gaussian random variables. Building on this, we then extend our analysis to establish novel concentration bounds for the SR. It is important to note that in this work, we define the SR as $\mu/\sigma^{2}$, where $\mu$ is the mean and $\sigma^{2}$ is the variance, deviating from the classical definition of $\mu/\sigma$.
@@
Let $X_{1},X_{2},\dots,X_{n}$ be i.i.d. sub-Gaussian random variables with mean $\mu=\mathbb{E}[X_{i}]$, variance $\sigma^{2}=\mathbb{E}[(X_{i}-\mu)^{2}]$, and sub-Gaussian parameter $v^{2}$, satisfying:
$\mathbb{E}[\exp(\lambda(X_{i}-\mu))]\leq\exp\left(\frac{\lambda^{2}v^{2}}{2}\right),\quad\forall\lambda\in\mathbb{R}.$
Then, for a constant $C>0$ we have:
$\mathbb{P}\left(|\bar{V}(n)-\sigma^{2}|>\epsilon\right)\leq 4\exp\left(-Cn\min\left(\frac{\epsilon^{2}}{v^{4}},\frac{\epsilon}{v^{2}}\right)\right).$
@@
See Appendix VII-F. ∎
Figure 1: Regret v/s time for UCB-VV, $\epsilon$-Greedy and VTS for (a) $\delta=0.1$ and (b) $\delta=0.5$ (c) UCB-VV v/s KL-UCB.
@@ IV-A Concentration inequality for Sharpe Ratio
The SR, defined as the ratio of expected reward to its standard deviation (or variance), is a standard metric for evaluating performance relative to risk in sequential decision-making frameworks. While our earlier research on bounded random variables derived path-dependent concentration inequalities for the SR [32], real-world scenarios often involve unbounded rewards. Extending results to the sub-Gaussian distributions not only broadens their applicability but also streamlines the analysis by removing dependencies of trajectory-specific bounds tied to bounded variables. Here, we establish a new concentration bound for the sample SR built on the variance bounds for sub-Gaussian random variables.
@@
Let $X_{1},X_{2},\dots,X_{n}$ be i.i.d. sub-Gaussian random variables with mean $\mu=\mathbb{E}[X_{i}]$, variance $\sigma^{2}=\mathbb{E}[(X_{i}-\mu)^{2}]$, and sub-Gaussian parameter $v^{2}$. If the true SR is $S=\frac{\mu}{\sigma^{2}}$ and empirical SR is $\bar{S}(n)=\frac{\bar{\mu}(n)}{\bar{V}(n)}$, there exists a constant $c>0$ depending on $\mu,\sigma^{2},v^{2}$ such that:
@@ IV-B UCB-Sharpe algorithm and its regret
Leveraging the above, we propose the UCB-Sharpe algorithm that selects arms with the high SR, balancing empirical performance with uncertainty in estimation. Let each arm $i\in\{1,\dots,K\}$ have unknown mean $\mu_{i}$, variance $\sigma_{i}^{2}$, and true SR $S_{i}=\mu_{i}/\sigma_{i}^{2}$. The UCB-Sharpe index at time $t$ is $\text{Index}_{i}(t)=\bar{S}_{i}(t)+\beta_{i}(t)$, where $\bar{S}_{i}(t)=\bar{\mu}_{i}/\bar{\sigma}_{i}^{2}$ is the estimated SR and $\beta_{i}(t)$ is the confidence bonus derived from the concentration inequality for the sample SR and is equal to $\beta_{i}(t)=\sqrt{\frac{1}{cs_{i}(t)}\log(4t^{2})}$. The steps are presented in Algorithm 3.
@@
3:        Compute sample mean $\bar{\mu}_{i}(t)$ and variance $\bar{V}_{(}t)$
4:        Compute estimated Sharpe ratio $\bar{S}_{i}(t)=\frac{\bar{\mu}_{i}}{\bar{\sigma}_{i}^{2}}$
@@
See Appendix VII-H. ∎
The regret scales logarithmically with $n$, similar to classical UCB algorithms, with constants depending on the sub-Gaussian parameter $v^{2}$, $\mu$ and $\sigma^{2}$.
@@ V-A Performance for UCB-VV
In Fig. 1, we compare the performance of our proposed algorithm UCB-VV with existing algorithms like VTS, $\epsilon$-greedy and KL-UCB. We run all algorithms for $2$ arms and a time horizon of $n=5000$. In Fig. 1a, for $\delta_{i}=0.1$, we see UCB-VV outperforms $\epsilon$-greedy for $\epsilon=0.1$ as well as $\epsilon=0.2$. The regret of VTS is the lowest. In Fig. 1b, for $\delta_{i}=0.2$, we see the regret decreases for all algorithms as the distinction between optimal and sub-optimal arms gets easier. However, the drop in regret is larger for UCB-VV and VTS. UCB-VV outperforms $\epsilon$-greedy for $\epsilon=0.1$ and $\epsilon=0.2$. The regret of VTS is again the lowest among all.
In case of $\delta_{i}=0.1$, we observe the regret for $\epsilon=0.1$ is larger than for $\epsilon=0.2$. This is because a larger $\epsilon$ implies larger exploration and thus the algorithm is able to pull each arm a greater number of times and can gather more data about all the arms, before converging on the optimal arm. But excessive exploration can lead to higher regret, as we can see in the case of $\delta_{i}=0.5$. We also observe that VTS, which is a variance-based Thompson sampling (TS) algorithm, outperforms our proposed algorithm for both sub-optimality gaps. While our proposed algorithm is optimal and offers theoretical guarantees, the superior performance of VTS highlights the need for a theoretical investigation of VTS. Thus providing the direction for future research, which could further enhance the applicability and impact of VTS.
KL-UCB employs KL divergence to construct tighter confidence intervals compared to traditional UCB methods, which rely on looser concentration inequalities such as Hoeffding’s or McDiarmid’s. It is asymptotically optimal for Bernoulli bandits, which also reflects in Figure 1c. To enable a fair comparison between UCB-VV and KL-UCB, we consider a two-armed ($K=2$) Bernoulli bandit setting. Each arm is defined by its success probability $p_{1}$ and $p_{2}$, respectively. The experiment ensures both algorithms face the same sub-optimal gap $\delta$ in expected reward. For UCB-VV, setting $p_{1}=0.5$ and $p_{2}=0.1838$ yields $\delta=0.1$, accounting for variance reward. For KL-UCB, the same gap is achieved with $p_{1}=0.5$ and $p_{2}=0.4$, as it focuses solely on mean reward. In both cases, the first arm is optimal. While KL-UCB selects arms based on maximum mean reward, UCB-VV selects based on maximum variance. This distinction is illustrated in Figure 1c, where both algorithms are evaluated under $\delta=0.1$ and $\delta=0.2$.
@@ V-B Performance for SHVV
We observe that as the number of arms increases, $e_{n}$ increases. We see in Fig. 2a2b, 2c, 2d and 2f, an error corresponding to the detection of optimal arm increases as $K$ jumps from $16-32$ and $32-64$ for a fixed budget of $2000$. The trend is similar for uniform sampling, although SHVV has the lowest error. This implies that as arms grow, the budget required is greater than provided $n$, for the algorithm to perform optimally with growing arms.
Also, we see in Fig. 2e, for SHVV, as the number of arms increases, $e_{n}$ decreases as the budget allocated is also increased with the increase in arms. The error of SHVV is the least compared to uniform sampling through all experiments.
@@ VI Case study: call option trading
Options are a widely used financial tool that investors employ to speculate on market movements or hedge against potential risks. The potential profits for option buyers are limitless, but their losses are capped, which is why they pay a premium. Conversely, option writers face potentially unlimited losses, while their profits are restricted, which is why they receive a premium. Call options are a fundamental derivative instrument in finance, granting the holder the right, but not the obligation, to purchase an underlying asset at the strike price, within a specified time period [9]. The buyer expects stock prices to rise, while the writer anticipates a decline. Volatility is a key driver of the option’s premium, as it reflects the market’s expectation of the asset’s future price fluctuations. The profit for the buyer is calculated as $\max(0,S_{t}-R)-P$, where $S_{t}$ is the stock price of the underlying asset at time $t$, $R$ is the strike price, i.e., the predetermined price at which the buyer can purchase the asset, and $P$ is the premium paid as a cost of purchasing the call option.

### v3

@@ II-E Lower bound on misallocation
Consider the MAB problem where the objective is to identify the arm with the highest variance. Let $\pi$ be an $\alpha$-consistent policy and $\Theta=\left\{\left\{\theta_{i}\right\}_{i=1}^{K}\colon\theta_{i}\in\mathcal{U}\right\}$ be the distribution model. Under Assumption 1, for any constant $C_{1}<1-\alpha$, $0<C_{2}<1$, with $f_{i}$ and $f_{*}$ as the reward distribution of the i-th arm and the optimal arm respectively, the model-specific misallocation satisfies,
$\liminf_{n\rightarrow\infty}\frac{\mathcal{M}_{\texttt{VV}}(n)}{\log n}\geq\sum_{i=1,i\neq*}^{K}\frac{C_{1}C_{2}}{I(f_{i},f_{*})}.$
@@
See Appendix VIII-C. ∎
From Theorem 1 and Theorem 2, we see UCB-VV achieves an upper bound on the misallocation of $\mathcal{O}\log(n)$ and matches a lower bound of the same order up to constant factors, demonstrating that our proposed algorithm is asymptotically optimal.
@@ III-A Upper bound on error for BAI
Parameter: $\bar{V}_{i}=0$
Initialize: $A_{1}\leftarrow K$, $r\leftarrow 1$
@@ III-B Lower bound on error for BAI
$G^{k}$ $=\begin{cases}\nu_{k}=\nu_{i}&i\neq k\\ \nu_{k}\sim U\left[\tilde{l}_{k},\tilde{u}_{k}\right]&i=k\end{cases}$
@@
$\mathbb{P}\left(\left|\frac{1}{n}\sum_{s=1}^{n}Z_{s}-\mathbb{E}[Z]\right|\geq\epsilon\right)\leq 2\exp\left(\frac{-2{n}\epsilon^{2}}{M^{2}}\right).$
@@ IV Sub-Gaussian Distributions and Sharpe Ratio
In this section, we extend our framework from bounded distributions to sub-Gaussian distributions. Specifically, we first derive a concentration inequality for the sample variance and standard deviation of sub-Gaussian random variables. Building on this, we extend our analysis to establish novel concentration bounds for the SR.
@@
Let $X_{1},X_{2},\dots,X_{n}$ be i.i.d. sub-Gaussian random variables with mean $\mu$, variance $\sigma^{2}$, and sub-Gaussian parameter $v^{2}$, satisfying
$\mathbb{E}\!\left[\exp\!\left(\lambda(X_{i}-\mu)\right)\right]\leq\exp\!\left(\frac{\lambda^{2}v^{2}}{2}\right),\qquad\forall\lambda\in\mathbb{R}$
Then, there exists a constant $C_{V}>0$ such that for any $\epsilon>0$,
$\mathbb{P}\left(|\bar{V}(n)-\sigma^{2}|>\epsilon\right)\leq 4\exp\!\left(-C_{V}n\min\!\left(\frac{\epsilon^{2}}{v^{4}},\frac{\epsilon}{v^{2}}\right)\!\right)$
@@
See Appendix VIII-F. ∎
Corollary 1 (Concentration of the Empirical Standard Deviation).
Under the assumptions of Theorem 5, we define the empirical standard deviation $\sqrt{\bar{V}(n)}$. Then for any $\epsilon>0$, there exists a constant $C>0$ such that
$\mathbb{P}\!\left(\left|\sqrt{\bar{V}(n)}-\sigma\right|>\epsilon\right)\leq 4\exp\!\left(-Cn\min\!\left(\frac{\epsilon^{2}}{v^{4}},\frac{\epsilon}{v^{2}}\right)\right).$
Figure 1: Misallocation v/s time for UCB-$V$, $\epsilon$-Greedy and VTS for (a) $\delta=0.1$, (b) $\delta=0.5$, and (c) $\delta=1$.
@@ IV-A Concentration inequality for Sharpe Ratio
The SR, defined as the ratio of expected reward to its standard deviation, is a standard metric for evaluating performance relative to risk in sequential decision-making frameworks. While our earlier research on bounded random variables derived path-dependent concentration inequalities for the SR [23], real-world scenarios often involve unbounded rewards. Extending results to the sub-Gaussian distributions not only broadens their applicability but also streamlines the analysis by removing dependencies of trajectory-specific bounds tied to bounded variables. Here, we establish a new concentration bound for the sample SR for sub-Gaussian random variables.
@@
Let $X_{1},X_{2},\dots,X_{n}$ be i.i.d. sub-Gaussian random variables with mean $\mu=\mathbb{E}[X_{i}]$, standard deviation $\sigma=\sqrt{\mathbb{E}[(X_{i}-\mu)^{2}]}$, and sub-Gaussian parameter $v^{2}$. If the true SR is $S=\frac{\mu}{\sigma}$ and empirical SR is $\bar{S}(n)=\frac{\bar{\mu}(n)}{\sqrt{\bar{V}(n)}}$, there exists a constant $c>0$ depending on $\mu,\sigma^{2},v^{2}$ such that:
@@ IV-B UCB-Sharpe algorithm and its regret
Leveraging the concentration inequality above, we propose the UCB-Sharpe algorithm that selects arms with the highest SR, balancing empirical performance with uncertainty in estimation. Let each arm $i\in\{1,\dots,K\}$ have unknown mean $\mu_{i}$, standard deviation $\sigma_{i}$, and true SR $S_{i}=\mu_{i}/\sigma_{i}$. The UCB-Sharpe index at time $t$ is $\text{Index}_{i}(t)=\bar{S}_{i}(t)+\beta_{i}(t)$, where $\bar{S}_{i}(t)=\bar{\mu}_{i}/\bar{\sigma}_{i}$ is the estimated SR and $\beta_{i}(t)$ is the confidence bonus derived from the concentration inequality for the sample SR and is equal to $\beta_{i}(t)=\sqrt{\frac{1}{Cs_{i}(t)}\log(4t^{2})}$. The steps are presented in Algorithm 3. .The constant $C$ required in the algorithm is a lower bound on the $c$ (from Theorem 6) for each of the $K$ arms.
@@
3:   Compute $\bar{\mu}_{i}(t)$ and $\sqrt{\bar{V}(t)}$
4:   Compute estimated SR $\bar{S}_{i}(t)=\frac{\bar{\mu}_{i}}{\bar{\sigma}_{i}}$
@@
9:end for
Algorithm 4 Variance Thompson Sampling - VTS [39]
Input: $x_{i,\max}(0)=0,s_{i}(0)=0$
1:     For each $t=1,2,\dots,K$ do
2:     Play arm $t$
3:     Update $(x_{i,\max}(t),s_{t}(t-1))$
4:    end For
5:     For each t=K+1,K+2,,nt=K+1,K+2,...,n do
6:     Sample $\tilde{\theta}_{i,t}\sim\mathrm{Pareto}\big(x_{i,\max,t-1},\,n_{i,t-1}\big)$
7:     Play arm $i(t)=\arg\max_{i\in[K]}\tilde{\theta}_{i,t}^{2}/12$ and observe reward $X_{i(t),t}$
8:     Update $(x_{i(t),\max}(t),s_{i(t)}(t-1))$
@@
See Appendix VIII-H. ∎
The regret scales logarithmically with $n$, similar to classical UCB algorithms, depending on the parameter $c>0$ which is a function of sub-Gaussian parameter $v^{2}$, $\mu$ and $\sigma^{2}$.
@@ V-A Performance for UCB-VV
In Fig. 1, we compare the performance of our proposed algorithm UCB-VV with existing algorithms like VTS and $\epsilon$-greedy. VTS [39] extends classical Thompson sampling (TS) to risk-aware decision making by incorporating uncertainty in variance rather than relying solely on mean rewards. While VTS [39] primarily considers Gaussian and Bernoulli reward models equipped with appropriate conjugate priors, our setting fundamentally differs in that, as rewards are bounded. For our numerical investigations, we model the reward of the $i$th arm as $X_{i}\sim\mathrm{Uniform}(0,\theta_{i})$, where the unknown parameter $\theta_{i}>0$ fully characterizes both mean and variance. Consequently, variance maximization is equivalent to the estimation and maximization of the scale parameter $\theta_{i}$. We adopt a non-informative prior $p(\theta_{i})\propto 1/\theta_{i}$, which is commonly used for scale parameters. Given $s_{i}(t)$ observations from arm $i$ up to time $t$, let $x_{i,\max}(t)=\max\left\{X_{i,1},\ldots,X_{i,s_{i}(t)}\right\}$ denote the sample maximum. Under the above prior, the posterior distribution of $\theta_{i}$ admits the closed-form Pareto density which exhibits a polynomially decaying heavy tail. In our implementation, TS for variance maximization proceeds by independently drawing $\tilde{\theta}i(t)\sim\mathrm{Pareto}\big(x_{i,\max}(t),s_{i}(t)\big)$ for each arm and selecting the arm that maximizes the sampled variance $\tilde{\theta}_{i}(t)^{2}/12$. The posterior parameters admit particularly simple online updates. After observing a new reward $X_{i,t}$ from arm $i$, the sufficient statistics are updated as
$s_{i}(t+1)=s_{i}(t)+1,\,\,x_{i,\max}(t+1)=\max\left\{x_{i,\max}(t),X_{i,t}\right\}.$
These two quantities fully characterize the posterior at every time step, making the proposed VTS implementation computationally lightweight.
We run all algorithms for $2$ arms and a time horizon of $n=5000$. In Fig. 1a, 1b and 1c we test our algorithm for both fixed values of ($\epsilon=0.1,0.2$) and a decaying schedule ($\epsilon_{t}=1/\sqrt{t}$). For UCB-VV, we use the theoretically derived confidence interval i.e, $C_{t,s_{i}}=\sqrt{\frac{2\log t}{s_{i}}}$. For $\delta_{i}=0.1$, we see UCB-VV outperforms $\epsilon$-greedy, but the misallocation of VTS is the lowest. In Fig. 1b, for $\delta_{i}=0.2$, we see the misallocation decreases for all algorithms as the distinction between optimal and sub-optimal arms gets easier. In case of $\delta_{i}=0.1$, we observe the misallocation for $\epsilon=0.1$ is larger than for $\epsilon=0.2$. This is because a larger $\epsilon$ implies larger exploration and thus the algorithm is able to pull each arm a greater number of times and can gather more data about all the arms, before converging on the optimal arm. But excessive exploration can lead to higher misallocation, as we can see in the case of $\delta_{i}=0.5$. However, the drop in misallocation is larger for UCB-VV and VTS, and the gap between the misallocation is also smaller. When the sub-optimal variance gap is moderate i.e. $\delta=0.1,0.5$, the problem is intrinsically difficult and requires sustained exploration of both arms. In this regime, the randomized exploration of VTS provides a more balanced trade-off between exploration and exploitation, while UCB-VV tends to be relatively conservative. As a result, VTS exhibits superior finite-time performance.
In contrast, when the sub-optimal gap is large i.e. $\delta=1$, as shown in Fig. 1c, the advantage shifts to UCB-VV. In this regime, the deterministic confidence bounds in UCB-VV separate the arms quickly, causing the algorithm to commit to the optimal arm after a short transient phase. The posterior of the uniform distribution under a non-informative prior is Pareto with heavy polynomial tails. Consequently, the VTS occasionally draws extremely large values for the suboptimal arm, even when it has already been sampled many times. These rare but persistent overestimates induce unnecessary exploration of the inferior arm. Although these events become less frequent over time, their persistence over a finite horizon is sufficient to give UCB-VV a clear performance advantage in the large-gap regime.
While our proposed algorithm is optimal and offers theoretical guarantees, the superior performance of VTS highlights the need for a theoretical investigation of VTS. Thus providing the direction for future research, which could further enhance the applicability and impact of VTS.
@@ V-B Performance for SHVV
We observe that as the number of arms increases, $e_{n}$ increases. We see in Fig. 2a, 2b, 2c, 2d and 2f, an error corresponding to the detection of optimal arm increases as $K$ jumps from 163216-32 and 326432-64 for a fixed budget of $2000$. The trend is similar for uniform sampling, although SHVV has the lowest error. This implies that as arms grow, the budget required is greater than provided $n$, for the algorithm to perform optimally with growing arms.
Also, we see in Fig. 2e, for SHVV, as the number of arms increases, $e_{n}$ decreases as the budget allocated is also increased proportionate to the increase in arms. The error of SHVV is the least compared to uniform sampling through all experiments.
@@ VI Case study: call option trading
Options are a widely used financial tool that investors employ to speculate on market movements or hedge against potential risks. The potential profits for option buyers are limitless, but their losses are capped, which is why they pay a premium. Conversely, option writers face potentially unlimited losses, while their profits are restricted, which is why they receive a premium. Call options are a fundamental derivative instrument in finance, granting the holder the right, but not the obligation, to purchase an underlying asset at the strike price, within a specified time period [29]. The buyer expects stock prices to rise, while the writer anticipates a decline. The call option payoff is given as $\max(0,S_{t}-R)$, where $S_{t}$ is the stock price of the underlying asset at time $t$, $R$ is the strike price, i.e., the predetermined price at which the buyer can purchase the asset. The profit for the buyer is calculated by subtracting the premium $P$ from the payoff. The premium paid is a cost of purchasing the call option.
VI-A Relation Between Variance and Call Option Payoff
In this paper, the primary theoretical objective is to study variance identification as a standalone online learning problem, in which the variance of each arm is treated as the reward signal. Accordingly, in the misallocation setting, the learner aims to identify the arm with the highest variance under bandit feedback. Although this departs from the classical mean-reward formulation, it remains a valid stochastic bandit setting: the learner sequentially interacts with unknown stochastic arms, observes noisy rewards, and seeks to minimize misallocation with respect to an unknown optimal arm. The call option trading example serves as a motivating application in which variance acts as a meaningful proxy for expected payoff due to the convex structure of the option payoff.
For a European call option, the payoff $f(S_{T})=\max(S_{T}-K,0)$ is a convex function of the terminal price $S_{T}$. Higher volatility increases the probability mass in the upper tail of the asset-price distribution, while the downside is capped at zero. As a result, the expected payoff of a call option increases with volatility. This fundamental property is well established in classical option pricing theory. In the Black-Scholes framework [5], the call price increases monotonically with the volatility parameter. Merton [29] further notes that a call option written on a more volatile stock is more valuable than one on a less volatile stock. A general economic justification is provided by the theory of mean-preserving spreads introduced by Rothschild and Stiglitz [32]. For random variables $X$ and $Y$, $X\leq_{r}Y\implies Y\overset{d}{=}X+Z,\mathbb{E}\left[Z\mid X\right]=0$, meaning that $Y$ is riskier than $X$. Their key result states that for any concave utility function $U$, $X\leq_{r}Y\implies\mathbb{E}\left[U(X)\right]\geq\mathbb{E}\left[U(Y)\right]$. The converse implication holds for convex functions. Since the European call payoff is convex in $S$. Hence, when volatility increases through a mean-preserving spread as in the Black-Scholes lognormal model, the expected call payoff necessarily increases with variance. This conclusion is also explicitly stated in standard texts. For example, Shreve [33] notes that as volatility increases, option prices increase in the Black-Scholes-Merton model.
To make the above convex-order and economic arguments fully explicit within a concrete asset-pricing model, we now establish the monotonic dependence of the European call price on the volatility parameter under the Black-Scholes-Merton framework. This result provides a direct analytical link between variance and call-option valuation and formally justifies the use of variance as a proxy objective in the present setting.
Lemma 6.
Under the Black-Scholes model with no dividends and constant risk-free rate $r$, the price of a European call option $C(t,x;\sigma)$ is a strictly increasing function of the volatility parameter $\sigma$. In particular, for fixed $(t,x,K,r,T)$,
$\sigma_{x}>\sigma_{y}\implies C\left(t,x;\sigma_{x}\right)>C\left(t,x;\sigma_{y}\right)$
Proof:
Let $\tau=T-t$ denote the time to maturity. Under the standard Black-Scholes assumptions of no dividends and a constant risk-free interest rate $r$, the price of a European call option at time $t$, with underlying price $S_{t}=x$, strike $K$, and volatility $\sigma$, is given by
$C(t,x;\sigma)=x\Phi(d_{1})-Ke^{-r\tau}\Phi(d_{2}),$
where
$d_{1}=\frac{\ln(x/K)+(r+\tfrac{1}{2}\sigma^{2})\tau}{\sigma\sqrt{\tau}},\qquad d_{2}=d_{1}-\sigma\sqrt{\tau},$
and $\Phi(\cdot)$ and $\phi(\cdot)$ denote the standard normal cumulative distribution function and density, respectively. To examine the sensitivity of the call price with respect to volatility, we compute the derivative $\partial C/\partial\sigma$, commonly referred to as the vega:
$\frac{\partial C}{\partial\sigma}=x\phi(d_{1})\frac{\partial d_{1}}{\partial\sigma}-Ke^{-r\tau}\phi(d_{2})\frac{\partial d_{2}}{\partial\sigma}.$
Let $A=\ln(x/K)+r\tau$. Then $d_{1}=\frac{A}{\sigma\sqrt{\tau}}+\tfrac{1}{2}\sigma\sqrt{\tau}$, which implies
$\frac{\partial d_{1}}{\partial\sigma}=-\frac{A}{\sigma^{2}\sqrt{\tau}}+\tfrac{1}{2}\sqrt{\tau},\qquad\frac{\partial d_{2}}{\partial\sigma}=-\frac{A}{\sigma^{2}\sqrt{\tau}}-\tfrac{1}{2}\sqrt{\tau}.$
Using the Black-Scholes identity $x\phi(d_{1})=Ke^{-r\tau}\phi(d_{2})$, we obtain
$\frac{\partial C}{\partial\sigma}=x\phi(d_{1})\left(\frac{\partial d_{1}}{\partial\sigma}-\frac{\partial d_{2}}{\partial\sigma}\right)=x\phi(d_{1})\sqrt{\tau}.$
Hence
$\frac{\partial C}{\partial\sigma}=x\sqrt{\tau}\phi(d_{1})>0.$
Since $x>0$, $\tau>0$, and $\phi(d_{1})>0$, the derivative is strictly positive for all admissible parameter values. Therefore, for fixed $(t,x,K,r)$ and $T$, the Black-Scholes call price is a strictly increasing function of the volatility $\sigma$. In particular, if $\sigma_{x}>\sigma_{y}$, then $C(t,x;\sigma_{x})>C(t,x;\sigma_{y})$, which provides a direct analytical confirmation of the general convex-order argument. ∎
It is important to clarify that our objective here is not to conflate variance with payoff. Rather, we employ variance as a proxy that captures the characteristics inherent to a call option’s utility without treating it as a direct optimization target. This case study serves as a qualitative illustration, demonstrating how variance-aware arm selection can be meaningfully integrated into a real-world decision-making pipeline. Crucially, we do not claim that UCB-VV or SHVV optimizes option profit; rather, they serve as a variance-driven screening mechanism to identify assets with high potential for favorable price excursions.